What volume of hydrogen will be produced at stp by the reaction 67.3 g of magnesium with excess water according to the following reaction?
1
Expert's answer
2020-11-23T06:50:14-0500

Mg + 2H2O "\rightarrow" Mg(OH)2 + H2

moles of Mg taken = (67.3/24) = 2.80 moles


Moles of H2 produce = 2.80moles


volume of H2 at STP = 2.80× 22.4 L = 62.8 L